# Manifolds, Random Matrices and Spectral Gaps: The geometric phases of generative diffusion

### v1

@@ Abstract
In this paper, we investigate the latent geometry of generative diffusion models under the manifold hypothesis. To this purpose, we analyze the spectrum of eigenvalues (and singular values) of the Jacobian of the score function, whose discontinuities (gaps) reveal the presence and dimensionality of distinct sub-manifolds. Using a statistical physics approach, we derive the spectral distributions and formulas for the spectral gaps under several distributional assumptions and we compare these theoretical predictions with the spectra estimated from trained networks. Our analysis reveals the existence of three distinct qualitative phases during the generative process: a trivial phase; a manifold coverage phase where the diffusion process fits the distribution internal to the manifold; a consolidation phase where the score becomes orthogonal to the manifold and all particles are projected on the support of the data. This ‘division of labor’ between different timescales provides an elegant explanation on why generative diffusion models are not affected by the manifold overfitting phenomenon that plagues likelihood-based models, since the internal distribution and the manifold geometry are produced at different time points during generation.
@@ 1 Introduction
Generative diffusion models have revolutionized the fields of computer vision and generative modeling, achieving state-of-the-art performance on image generation (Ho et al.,, 2020; Song and Ermon,, 2019; S. et al.,, 2021) and video generation (Ho et al.,, 2022; Singer et al.,, 2022; Blattmann et al.,, 2023; B. et al., 2024b, ). Generative diffusion models synthesize images through a stochastic dynamical denoising process. Experimental and theoretical arguments suggest that different features such as frequency modes and class labels are generated at different times during the process. For example, it has been shown that separation between isolated classes, like in the case of mixture of Gaussian models, happens at critical phase transition points of spontaneous symmetry breaking (speciation events) (B. et al., 2024a, ). It is also well known that subspaces corresponding to different frequency models emerge at different times of diffusion (Kingma and Gao,, 2024). This idea has been recently refined by (Kadkhodaie et al.,, 2024), who showed that diffusion models give rise to a local decomposition of the image manifold into a basis of geometry-adaptive harmonic basis functions. These decomposition phenomena cannot be directly explained in terms of critical phase transitions as they are fundamentally linear processes. In this paper, we will provide a precise theoretical analysis of the separation of subspaces for data defined on low dimensional linear manifolds.
Our main contributions are: I) an in-depth theoretical random-matrix analysis of the distribution of Jacobian spectra in diffusion models on linear manifolds and II) a detailed experimental analysis of Jacobian spectra extracted from trained networks on linear manifolds and on image datasets. The analysis of this spectra is important as it provides a detailed picture of the latent geometry that guides the generative diffusion process. We show that the linear theory predicts several phenomena that we observed in trained networks. Based on our result, we divide the generative process in three qualitatively different phases: trivial phase, manifold coverage phase and manifold consolidation phase. Using these concepts, we provide an elegant explanation of why diffusion models can avoid the manifold overfitting pathology that characterizes likelihood-based models (Loaiza-Ganem et al.,, 2022).
@@ 2 The manifold hypothesis
The manifold hypothesis states that the distribution on natural data, such as images and sound recordings, is (approximately) supported on a $m-$dimensional manifolds $\mathcal{M}$ embedded in a larger euclidean ambient space $\mathds{R}^{d}$ (Peyré,, 2009; Fefferman et al.,, 2016). While a data probability distribution supported on a $m<d$ manifold $\mathcal{M}$ cannot be expressed using a proper density function, we loosely define such density as
$p_{0}(\boldsymbol{x})=\delta_{\mathcal{M}}(\boldsymbol{x})\,\rho_{\text{int}}(\boldsymbol{x})~{},$ (1)
@@
Figure 1: Visualization of the gaps in the spectrum of the (negative) Jacobian of the score for data supported on a latent manifold. Blue line: idealize spectrum of distribution with uniform internal density; Orange line: spectrum of a more realistic distribution.
2.1 Manifold overfitting in likelihood-based models
Likelihood-based generative models are defined by a highly parameterized likelihood function $f(\boldsymbol{x};\boldsymbol{\theta})$, whose parameters are trained by minimizing the loss
which maximizes the probability of the data given the model. This maximum likelihood loss is minimized if $f(\boldsymbol{x};\boldsymbol{\theta})=p_{0}(\boldsymbol{x})$. Unfortunately, the divergence of $p_{0}(\boldsymbol{x})$ on the manifold implies that the model density $f(\boldsymbol{x};\boldsymbol{\theta})$ on the manifold becomes larger and larger during training. More problematically, the optimization problem becomes almost insensitive to the internal density $\rho_{\text{int}}(\boldsymbol{x})$. This phenomenon is called manifold overfitting (Loaiza-Ganem et al.,, 2022), since the trained model fits the manifold while ignoring its internal density, resulting in poor generation. While generative diffusion models have often been characterized in terms of a likelihood function, experimental evidence suggests that they are not affected by this pathology. In the rest of the paper, we will provide an elegant explanation of their behavior by analyzing their latent geometry.
@@ 3 Background on generative diffusion models
where $\text{d}\boldsymbol{Z}_{t}$ is a standard Brownian motion. The formal solution of Eq. (3) can be given in term of the heat kernel: $p(\boldsymbol{x}_{t},t)=\mathbb{E}_{\boldsymbol{x}_{0}\sim p_{0}}\!\left[\frac{1}{\sqrt{2\pi t}}e^{-\frac{\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{0}\right\rVert_{2}^{2}}{2t}}\right]~{}.$ The target distribution $p_{0}(\boldsymbol{x})$ is then recovered by reversing the diffusion process (Anderson,, 1982). We initialize this reverse process from $\boldsymbol{x}_{t_{f}}\sim\mathcal{N}\left(0,t_{f}I_{d}\right)$ at some large time $t_{f}$, which evolves backward according to
@@
The function $s(\boldsymbol{x},t)=\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ is is the so-called score function. From a set of training points $\{\boldsymbol{y}^{1},\dots,\boldsymbol{y}^{N}\}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}p_{0}$, we can train a neural approximation of $s(\boldsymbol{x},t)$ by learning a denoising autoencoder $\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}(\mathbf{x},t)$, which is trained to recover the standardized noise $\epsilon_{t}$ from the noisy state $x_{t}=x_{0}+\sqrt{t}\epsilon_{t}$ (Hyvärinen and Dayan,, 2005; Vincent,, 2011; Ho et al.,, 2020). In order to avoid singularities in neural network output, the learned score is parametrized as $\hat{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)=-\frac{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}(\boldsymbol{x},t)}{\sqrt{t}}~{}.$
It is also convenient to define the support score $\tilde{s}_{\mathcal{M}}(\boldsymbol{x},t)$, defined as the score function obtained from the uniform data distribution $\tilde{p}_{0}(\boldsymbol{x})=\frac{1}{|\mathcal{M}|}\delta_{\mathcal{M}}(\boldsymbol{x})$.
@@ 4 Dynamic latent manifolds and spectral gaps
Consider a generative diffusion model with $p_{0}(\boldsymbol{x})$ defined on a $d$-dimensional manifold $\mathcal{M}_{0}$ according to Eq. (2). In the course of the diffusion process, we can define a time-dependent family of stable latent manifolds
@@
where the negative semi-definiteness (n.s.d.) is a stability condition on the Jacobian matrix $J_{\mathcal{M}}(\boldsymbol{x},t)$ of the support score $\tilde{s}_{\mathcal{M}}(\boldsymbol{x}^{*},t)$. Due to the noise, the diffusing particles are likely to be found in shells of radius $\sqrt{t}$ around each point of the latent manifold.
For a small perturbation $\boldsymbol{p}$ around a point $\boldsymbol{x^{*}}$ on the latent manifold at time $t$, the score function is well approximated by its linearization:
$s(\boldsymbol{p},t)\approx J(\boldsymbol{x}^{*},t)~{}\boldsymbol{p}=-\sum_{j}\left(\boldsymbol{v}_{j}\cdot\boldsymbol{p}\right)\lambda_{j}(\boldsymbol{x}^{*},t)\boldsymbol{v}_{j}~{}~{},$ (6)
where $J(\boldsymbol{x}^{*},t)$ is the Jacobian of the score and the $\boldsymbol{v}_{j}$ and $\lambda_{j}(\boldsymbol{x}^{*},t)$ are respectively the $j$-th eigenvector and the associated eigenvalue of $-J(\boldsymbol{x}^{*},t)$. The spectrum of eigenvalues provides detailed information concerning the local geometry of the manifold around. Perturbations aligned to the tangent space of $\mathcal{M}_{t}$ correspond to small eigenvalues, while orthogonal perturbations correspond to high eigenvalues, as the score tends to push the stochastic dynamics towards its fixed-points. Therefore, we can estimate the dimensionality of the manifold from the location of a drop (i.e. a sharp change) in the sorted spectrum of eigenvalues (Stanczuk et al.,, 2022). This is visualized in Fig. 1. This drop corresponds exactly to a gap (i.e. a separation) in the eigenvalues spectrum; in the following we will refer to both as gaps.
@@
Consider the situation where the internal density $\rho_{\text{int}}(\boldsymbol{x})$ is not locally flat around a point $\boldsymbol{x}^{*}\in\mathcal{M}_{0}$. In this case, at a finite time $t$ the actual score function does not vanish on the latent manifold $\mathcal{M}_{t}$ as there is a gradient of the log-density along the tangent directions. This implies that the spectrum of tangent eigenvalues can have a series of sub-gaps with separate different tangent subspaces with different ‘local variance’. In image generation tasks, these subspaces are often associated with different frequency modes, as noted in (Kingma and Gao,, 2024).
Consequently, we can quantify the sensitivity to the internal density at time $t$ by studying the statistics and temporal evolution of intermediate gaps
@@ 5 Phenomenology of generative diffusion on manifolds
This section contains an intuitive picture that follows from our theoretical results on linear models, which we will fully outline in the next section. The theory considers the case of a linear manifold with Gaussian internal distribution. While only linear models are theoretically tractable, we conjecture that their phenomenology captures the main features of subspace separation in the tangent space of curved manifolds. We validated the theory using networks trained on both linear data and highly non-linear data such as natural images (see section 8). Based on the dynamics of the spectral gaps, we found that the generative dynamics of $\boldsymbol{x}_{t}$ according to Eq. (4) can be separated into three distinct phases. The phase separation do not correspond to singularities as there are cross-over events, not genuine phase transitions. During all our analysis we will exclusively work with eigenvalues since, in the linear manifold model, the Jacobian of the true score is symmetric. The same phenomenology is nevertheless fully appreaciable when using the singular values in our experiments with neural approximations of the score.
@@ 5.1 Phase I: The trivial phase
In the trivial phase, the diffusing particle moves according to the noise distribution without strong biases towards the manifold directions. In this dynamic regime, the latent manifold $\mathcal{M}_{t}$ is a single point surrounded by an isotropic quadratic well of potential. The spectral gaps are not visible and all eigenvalues have approximately the same value due to the isotropy of the noise distribution. This trivial phase is analogous to the initial phases described in (Raya and Ambrogioni,, 2024) and (B. et al., 2024a, ).
@@ 5.2 Phase II: Manifold coverage
The manifold coverage phase begins with the opening of the first of a series of spectral gaps corresponding to local subspaces. In this phase, different subspaces with different variances can therefore be identified by intermediate gaps in the spectra, as sketched in Fig. 2. When the intermediate gaps are opened, the diffusing particles spread across the relevant manifold directions according to their relative variances. In other words, during this regime of generative diffusion, the process fits the distribution of the data internal to the manifold.
We assume low-rank covariance $\Sigma=FF^{\top}$ for the data distribution. In terms of random matrix theory, the gap-forming phenomenology has two distinct processes: the emergence of intermediate gaps (i.e. steps in the dimensionality plot) between separated bulks of the spectrum, the opening of a final gap that allows to infer the dimensionality of the full manifold. Our analysis gives us the time scale at which such intermediate gaps are maximally opened, i.e.
@@
where $\gamma_{-}(\sigma_{k+1})$ and $\gamma_{+}(\sigma_{k})$ are specific eigenvalues of $\Sigma$ (see Fig. 9) associated to two hierarchically consecutive variances (see Appendix A.2 for an exhaustive analysis). In most cases, when $\sigma_{k+1}^{2}\ll\sigma_{k}^{2}$, the dependence on the two variances is $\mathcal{O}(\sigma_{k}\cdot\sigma_{k+1})$. This is the timescale where the score is maximally sensitive to the relative variance of the two subspaces, which guides the particles toward the correct internal distribution.

### v4

@@ Abstract
In this paper, we investigate the latent geometry of generative diffusion models under the manifold hypothesis. For this purpose, we analyze the spectrum of eigenvalues (and singular values) of the Jacobian of the score function, whose discontinuities (gaps) reveal the presence and dimensionality of distinct sub-manifolds. Using a statistical physics approach, we derive the spectral distributions and formulas for the spectral gaps under several distributional assumptions, and we compare these theoretical predictions with the spectra estimated from trained networks. Our analysis reveals the existence of three distinct qualitative phases during the generative process: a trivial phase; a manifold coverage phase where the diffusion process fits the distribution internal to the manifold; a consolidation phase where the score becomes orthogonal to the manifold and all particles are projected on the support of the data. This ‘division of labor’ between different timescales provides an elegant explanation of why generative diffusion models are not affected by the manifold overfitting phenomenon that plagues likelihood-based models, since the internal distribution and the manifold geometry are produced at different time points during generation.
@@ 1 Introduction
Generative diffusion models have revolutionized the fields of computer vision and generative modeling, achieving state-of-the-art performance on image generation (Ho et al.,, 2020; Song and Ermon,, 2019; Song et al.,, 2021) and video generation (Ho et al.,, 2022; Singer et al.,, 2022; Blattmann et al.,, 2023; Brooks et al.,, 2024). Generative diffusion models synthesize images through a stochastic dynamical denoising process. Experimental and theoretical arguments suggest that different features such as frequency modes and class labels are generated at different times during the process. For example, it has been shown that separation between isolated classes, as in the case of mixture of Gaussian models, happens at critical phase transition points of spontaneous symmetry breaking (speciation events) (Biroli et al.,, 2024). It is also well known that subspaces corresponding to different frequency modes emerge at different times of diffusion (Kingma and Gao,, 2024). This idea has been recently refined by (Kadkhodaie et al.,, 2024), who showed that diffusion models give rise to a local decomposition of the image manifold into a basis of geometry-adaptive harmonic basis functions. These decomposition phenomena cannot be directly explained in terms of critical phase transitions as they are fundamentally linear processes. In this paper, we will provide a precise theoretical analysis of the separation of subspaces for data defined on low dimensional linear manifolds.
Our main contributions are: I) an in-depth theoretical random-matrix analysis of the distribution of Jacobian spectra in diffusion models on linear manifolds and II) a detailed experimental analysis of Jacobian spectra extracted from trained networks on linear manifolds and on image datasets. The analysis of these spectra is important as it provides a detailed picture of the latent geometry that guides the generative diffusion process. We show that the linear theory predicts several phenomena that we observed in trained networks. Based on our result, we divide the generative process into three qualitatively different phases: trivial phase, manifold coverage phase and manifold consolidation phase. Using these concepts, we provide a concise explanation of why diffusion models can avoid the manifold overfitting pathology that characterizes likelihood-based generative models (Loaiza-Ganem et al.,, 2022).
@@ 2 The manifold hypothesis
The manifold hypothesis states that the distribution on natural data, such as images and sound recordings, is (approximately) supported on a $m-$dimensional manifold $\mathcal{M}$ embedded in a larger euclidean ambient space $\mathds{R}^{d}$ (Peyré,, 2009; Fefferman et al.,, 2016). While a data probability distribution supported on a $m<d$ manifold $\mathcal{M}$ cannot be expressed using a proper density function, we loosely define such density as
$p_{0}(\boldsymbol{x})=\delta_{\mathcal{M}}(\boldsymbol{x})\,\rho(\boldsymbol{x})~{},$ (1)
@@
$\boldsymbol{y}^{\mu}=\Phi\left(F\boldsymbol{z}^{\mu}\right),$ (2)
where $\Phi(\boldsymbol{x})$ is a non-linear function of the $d$-dimensional vector $\boldsymbol{x}$, $F\in\mathbb{R}^{d\times m}$ is a rectangular matrix that projects latent vectors $\boldsymbol{z}^{\mu}$ on the manifold. In this work we are going to study generative diffusion while aligning with this modeling frame.
Figure 1: (a) Visualization of the gaps in the spectrum of the (negative) Jacobian of the score for data supported on a latent manifold. Blue line: idealized spectrum of distribution with uniform internal density; Orange line: spectrum of a more realistic distribution. (b) Sketch of the local structure of data-manifold with tangent and orthogonal components of the score function. (c) Sketch of the geometric phases of generative diffusion and their trace measurable from the eigenspectrum.
@@ 3 Background on generative diffusion models
where $\text{d}\boldsymbol{Z}_{t}$ is a standard Brownian motion. The formal solution of Eq. (3) can be given in terms of the heat kernel $p(\boldsymbol{x}_{t},t)=\mathbb{E}_{\boldsymbol{x}_{0}\sim p_{0}}\!\left[\frac{1}{\sqrt{2\pi t}}e^{-\frac{\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{0}\right\rVert_{2}^{2}}{2t}}\right].$ The target distribution $p_{0}(\boldsymbol{x})$ is then recovered by reversing the diffusion process (Anderson,, 1982). We initialize this reverse process from $\boldsymbol{x}_{t_{f}}\sim\mathcal{N}\left(0,t_{f}I_{d}\right)$ at some large time $t_{f}$, which evolves backward in time according to
@@
The function $s(\boldsymbol{x},t)=\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x})$ is the so-called score function. From a set of training points $\{\boldsymbol{y}^{1},\dots,\boldsymbol{y}^{N}\}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}p_{0}$, we can train a neural approximation of $s(\boldsymbol{x},t)$ by learning a denoising autoencoder $\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}(\mathbf{x},t)$, which is trained to recover the standardized noise $\epsilon_{t}$ from the noisy state $x_{t}=x_{0}+\sqrt{t}\epsilon_{t}$ (Hyvärinen and Dayan,, 2005; Vincent,, 2011; Ho et al.,, 2020). In order to avoid singularities in neural network output, the learned score is parametrized as $\hat{s}_{\boldsymbol{\theta}}(\boldsymbol{x},t)=-\frac{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{\theta}}(\boldsymbol{x},t)}{\sqrt{t}}~{}.$
@@ 4 Dynamic latent manifolds and spectral gaps
Consider a generative diffusion model with $p_{0}(\boldsymbol{x})$ defined on a $d$-dimensional manifold $\mathcal{M}$ according to Eq. (1). In the course of the diffusion process, we can define a time-dependent locus of points
@@
that we name stable latent set of the process. In Eq. 5 we have used $\mathcal{M}\equiv\mathcal{M}_{0}$. The negative semi-definiteness (n.s.d.) is a stability condition on the Jacobian matrix $J_{\mathcal{M}}(\boldsymbol{x},t)$ of the support score $\tilde{s}_{\mathcal{M}}(\boldsymbol{x}^{*},t)$, defined as the score function obtained from the uniform data distribution $\tilde{p}_{0}(\boldsymbol{x})=\frac{1}{|\mathcal{M}|}\delta_{\mathcal{M}}(\boldsymbol{x})$. Due to the noise, the diffusing particles typically explore shells of a radius that concentrates on $\sqrt{t}$ around each point of the latent stable set. For a small perturbation $\boldsymbol{p}$ around a point $\boldsymbol{x^{*}}$ on the latent manifold at time $t$, the score function is well approximated by its linearization:
$s(\boldsymbol{p},t)\approx J(\boldsymbol{x}^{*},t)~{}\boldsymbol{p}=-\sum\nolimits_{j}\left(\boldsymbol{v}_{j}\cdot\boldsymbol{p}\right)\lambda_{j}(\boldsymbol{x}^{*},t)\boldsymbol{v}_{j}~{},$ (6)
where $J(\boldsymbol{x}^{*},t)$ is the Jacobian of the score and the $\boldsymbol{v}_{j}$ and $\lambda_{j}(\boldsymbol{x}^{*},t)$ are respectively the $j$-th eigenvector and the associated eigenvalue of $-J(\boldsymbol{x}^{*},t)$. The spectrum of eigenvalues provides detailed information concerning the local geometry of stable latent set. Perturbations aligned with the tangent space of $\mathcal{M}_{t}$ correspond to small eigenvalues, while orthogonal perturbations correspond to high eigenvalues, as the score tends to push the stochastic dynamics towards its fixed-points. Therefore, we can estimate the dimensionality of the manifold from the location of a drop (i.e. a sharp change) in the sorted spectrum of eigenvalues (Stanczuk et al.,, 2022). This is visualized in Fig. 1, panels (a) and (b). This drop corresponds exactly to a gap (i.e. a separation) in the eigenvalues spectrum; in the following, we will refer to both as gaps.
@@ 4.1 Subspaces and intermediate gaps
Consider the situation where the internal density $\rho_{\text{int}}(\boldsymbol{x})$ is not locally flat around a point $\boldsymbol{x}^{*}\in\mathcal{M}$. In this case, at a finite time $t$ the actual score function does not vanish on the latent stable set $\mathcal{M}_{t}$ as there is a gradient of the log-density along the tangent directions. This implies that the spectrum of tangent eigenvalues can have a series of sub-gaps with separate different tangent subspaces with different ‘local variance’. In image generation tasks, these subspaces are often associated with different frequency modes, as noted in (Kingma and Gao,, 2024). Consequently, we can quantify the sensitivity to the internal density at time $t$ by studying the statistics and temporal evolution of intermediate gaps
@@ 5 Phenomenology of generative diffusion on manifolds
This section contains an intuitive picture that follows from our theoretical results on linear models, which we will fully outline in the next section. The theory considers the case of a linear manifold with Gaussian internal distribution. A linear-manifold data-model is made of a set of points
$\boldsymbol{y}^{\mu}=F\boldsymbol{z}^{\mu},$ (8)
where $F$ and $\boldsymbol{z}^{\mu}$ have been introduced in Section 2. While only linear models are theoretically tractable, we conjecture that their phenomenology captures the main features of subspace separation in the tangent space of curved manifolds (see Supp. D for further details). We validated the theory using networks trained on both linear data and highly non-linear data such as natural images (see Sections 7 and 8). Based on the dynamics of the spectral gaps, we found that the generative dynamics of $\boldsymbol{x}_{t}$ according to Eq. (4) can be separated into three distinct phases. The phase separation does not correspond to singularities as there are cross-over events, not genuine phase transitions. During all our analysis we will exclusively work with eigenvalues since, in the linear manifold model, the Jacobian of the true score is symmetric. The same phenomenology is nevertheless fully appreciable when using the singular values in our experiments with neural approximations of the score.
@@ 5.1 Phase I: The trivial phase
In the trivial phase, the diffusing particle moves according to the noise distribution without strong biases towards the manifold directions. In this dynamic regime, the stable latent set $\mathcal{M}_{t}$ is a single point surrounded by an isotropic quadratic well of potential. The spectral gaps are not visible and all eigenvalues have approximately the same value due to the isotropy of the noise distribution. This trivial phase is analogous to the initial phases described in (Raya and Ambrogioni,, 2024) and (Biroli et al.,, 2024).
@@ 5.2 Phase II: Manifold coverage
The manifold coverage phase begins with the opening of the first of a series of spectral gaps corresponding to local subspaces. In this phase, different subspaces with different variances can therefore be identified by intermediate gaps in the spectra, as sketched in Fig. 1, panel (c). When the intermediate gaps are opened, the diffusing particles spread across the manifold directions according to their relative variances. In other words, during this regime of generative diffusion, the process fits the distribution of the data internal to the manifold.
We assume low-rank covariance $\Sigma=FF^{\top}$ for the data distribution. In terms of random matrix theory, the gap-forming phenomenology has two distinct processes: the emergence of intermediate gaps (i.e. steps in the dimensionality plot) between separated bulks of the spectrum, and the opening of a final gap that allows to infer the dimensionality of the full manifold. Our analysis gives us the time scale at which such intermediate gaps are maximally opened, i.e.
@@
where $\gamma_{-}(\sigma_{k+1})$ and $\gamma_{+}(\sigma_{k})$ are specific eigenvalues of $\Sigma$ (see Fig. 8) associated with two hierarchically consecutive variances (see Supp. A.2 for an exhaustive analysis). In most cases, when $\sigma_{k+1}^{2}\ll\sigma_{k}^{2}$, the dependence on the two variances is $\mathcal{O}(\sigma_{k}\cdot\sigma_{k+1})$. This is the timescale where the score is maximally sensitive to the relative variance of the two subspaces, which guides the particles toward the correct internal distribution.